Note: Students will solve this section and return it to teacher for checking.
1. Centre of gravity of rectangle is ___________
(a) it's center. (b) intersection of diagonals. (c) middle point of axis
2. A couple is formed by two _____ forces.
(a) like. (b)unlike. (c) Different
3. Centre of gravity of meter rod is ________
(a) center (b) intersection of diagonal. (c) intersection of medians
4. The torque produced in a body due to couple is equal to the product of one of the
force and________
(a)Couple. (b)Force. (c) Couple arm
5.When an object is resting in smooth horizontal surface it's weight is balanced by__
(a)Torque. (b)Reactional force (c) Couple
Q2: Write difference between couple and torque?
Q3: Find moment of couple produce by force of 50 N acting on two sides of stearing
wheel having radius of 20cm.
1
Expert's answer
2020-04-23T12:52:48-0400

1)B; 2)B; 3)A; 4)C; 5)B 7)"M=F\times2r=50N\times0.4m=20N\times m"

6) "\overline{M} =\overline{r}\times\overline{F}" couple

"\overline{-M'} =\overline{r}\times\overline{F}torque\\7)M=F\times2r=50N\times0.4m=20N\times m"



6)